OpenStudy (anonymous):

MEDAL AWARDED FOR HELP (Polynomials) Bernard is asked to create a swimming pool and deck design for a new series of model homes currently being designed. Each home will have varying pool sizes depending on the lot size; however, there will be some similarities, so polynomial expressions will be used. To fit the dimensions of the various land spaces, he realizes that the length of the pool has to be three times the size of the width plus 3 and the depth has to be 7 less than twice the length. Task 1 Create expressions to represent the dimensions of the pool

OpenStudy (anonymous):

I have this - Width = W, L = 3W + 3, D = 7 – 2L Is that correct

OpenStudy (anonymous):

I'm realllllly desperate. This is only one of more questions on this and I need help. If someone helps me out they can have my kik or something

OpenStudy (anonymous):

Also next question is : Create a variable to represent the width of the pool. Determine the expressions to represent length and the depth in terms of the width variable.

OpenStudy (anonymous):

Yes that is correct

OpenStudy (anonymous):

Thank you @Matthew071 can you continue helping? :)))))

OpenStudy (anonymous):

Sure :)

OpenStudy (anonymous):

What is your next question

OpenStudy (anonymous):

Create a variable to represent the width of the pool. Determine the expressions to represent length and the depth in terms of the width variable.

OpenStudy (anonymous):

Didnt you just already do that? First I think its just asking is just create a variable for width. A variable can be any letter or symbol. So what you did was create Width=W and W is the variable of width. Now in terms of width means just put the equations as in terms of width. So as you did L=3W or Length is 3 times the width plus 3.

OpenStudy (anonymous):

2. Determine the polynomial expression to represent the area of land space that the pool will cover. I'm sorry i copied the wrong one

OpenStudy (anonymous):

Its ok now lemme see

OpenStudy (anonymous):

Ok so we assume the pool is a rectangle so all we need to know is that Area=LeangthxWidth

OpenStudy (anonymous):

So now we just plug in that formula where Area=(W)(3W+3)

OpenStudy (anonymous):

Then we just simplify the equation where Area=3W^2+3W

OpenStudy (anonymous):

And that will be all

OpenStudy (anonymous):

I see. Thank you. More questions coming. I don't know how to thank you more than just a medal. Any ideas?

OpenStudy (anonymous):

3. Using the area and the depth expressions, determine the polynomial expression that is used to represent the volume also heres another one

OpenStudy (anonymous):

Keep the questions coming

OpenStudy (anonymous):

Ok now you know that Area=3W^2+3W now all you have to do is Plug into the formula for Volume. V=LWH or V=(Area)(D), Having said that we just plug in the variables and we have V=(3W^2+3W)(7-2L)

OpenStudy (anonymous):

Oh by the way I forgot to mention on the first part of the problem, since Depth is in terms of L we have to make it in terms of W

OpenStudy (anonymous):

So would it be 6w -4

OpenStudy (anonymous):

My Bad Just a simple fix D=7-2(3W) D=7-6W V=(3W^2+3W)(7-6W) Now do the FOIL method and you get Volume=-18w^3-3W^2+21W

OpenStudy (anonymous):

My bad again

OpenStudy (anonymous):

Volume=(3W^2+3W)(1-6W)

OpenStudy (anonymous):

you were correct on one part of the depth equation except for one part

OpenStudy (anonymous):

Okay thank you. You're a godsend. Hold on i'm looking at the assignment

OpenStudy (anonymous):

Ok and your welcome

OpenStudy (anonymous):

4. Classify the volume polynomial by degree and number of terms. It is a trinomial correct?

OpenStudy (anonymous):

Indeed and the degree is the term with the highest power in the equation so what is the highest power?

OpenStudy (anonymous):

So we know the highest power is cubed so what does that tell you?

OpenStudy (anonymous):

As in \[x^3+x^2+x\] the degree is 3 because the highest power is the degree

OpenStudy (anonymous):

Catch my drift?

OpenStudy (anonymous):

I undestand. its a 3rd degree trinomial

OpenStudy (anonymous):

CORRECTAMUNDO!!!

OpenStudy (anonymous):

:D

OpenStudy (anonymous):

Next i have to draw a diagram and i dont think you can help me with that, so thank you so much for your help. i wanna thank you more than just a medal.... How can I?

OpenStudy (anonymous):

Anything really :)

OpenStudy (anonymous):

Cool, I'll kik you. Thank you friend

OpenStudy (anonymous):

No problamo